# Towards a robust model-independent test of the DAMA/LIBRA dark matter signal: ANAIS–112 results with six years of data

### v1

@@ Introduction
While astronomical and cosmological evidence about dark matter (DM) accumulates, its nature remains elusive [1]. Over the past thirty years, various excesses above background have been reported in both indirect and direct dark matter searches, most of which have been resolved. However, a few of them persist and deserve further investigation [2]. One of the most puzzling results is that from the DAMA/LIBRA experiment, taking data at the Laboratori Nazionali del Gran Sasso (Italy), which reports a positive detection for more than twenty years [3, 4]. DAMA/LIBRA detectors observe an annual modulation in the detection rate at low energies, below 6 keV (in electron-equivalent energy 111In this article, we will use keV for referring to electron-equivalent energies while nuclear recoils energies will be expressed in keVnr.), consistent with the predictions of weakly interacting massive particles (WIMPs) distributed in the galactic halo [6, 7], and having cross-sections with nucleons in the range of cm2 [3]. The statistical significance of such a detection is overwhelming, but the origin of such a modulation has not yet been determined. No other experiment has observed a compatible signal [8, 9], but the uncertainties and unknowns in both the dark matter particle and halo models make it very difficult to disprove the interpretation of the result in terms of a DM signal in a model-independent way [2]. This is the objective of ongoing or recently decommissioned experiments, such as ANAIS–112 [10, 11, 12, 13, 14] and COSINE–100 [15, 16], or other upcoming projects like COSINE-100U [17], COSINUS [18], SABRE [19, 20, 21], and PICOLON [22].
@@
The DAMA/LIBRA experiment is expected to have completed the data taking by the end of 2024. The final results should then be released. The COSINE–100 experiment was in operation from 2016 to 2023 in the YangYang Underground Laboratory in Korea. After concluding data collection, the COSINE–100 collaboration recently released results from the full dataset [23], corresponding to 6.4 years of operation and a total effective exposure of 358 kgyr. The ANAIS–112 experiment, in operation since August 2017 at the Canfranc Underground Laboratory (LSC) in Spain, consists of 112.5 kg of NaI(Tl), distributed in nine detectors, labeled in the following D0 to D8, and produced by the same provider as those used in the COSINE–100 experiment (Alpha Spectra Inc.). A blank module, similar in design to the other modules but without the NaI(Tl) crystal, operates in the same experimental space within an independent shielding, but integrated into the same data acquisition system, to monitor non-NaI(Tl) scintillation events. Details on the ANAIS–112 experimental set-up, detector performance and previous results can be found elsewhere [10, 11, 12, 13, 24, 14]. ANAIS–112 will complete more than 8 years of data by the end of 2025, achieving a sensitivity exceeding the 5 confidence level according to our prospects. Here, we present the analysis of the annual modulation corresponding to six years of ANAIS–112 data for a total effective exposure of 625.75 kgyr, which confirms our sensitivity estimates.
@@
First, we briefly describe the key features of the analysis pipeline applied to the 6 years data, which in some aspects differ from that applied to obtain previous results [11, 12, 13, 14].
@@
A new filtering procedure to remove non-bulk scintillation events is applied, following Ref. [24, 26], as done in Ref. [14] for the reanalysis of the 3-year exposure. This filtering protocol is based on Boosted Decision Trees (BDT) and the training does not use background events at all, but uses events from 252Cf calibrations in the [1–2] keV energy region resulting mainly from nuclear elastic scattering from neutrons as signal and non-NaI(Tl) bulk scintillation events from the blank module in an equivalent energy range, as noise. The efficiency of this filtering procedure is derived both from 252Cf and 109Cd calibrations data, and their respective time evolution in the [1–6] keV energy region along the data taking is shown in the Figure 8 of the Supplemental Material [25]. By considering the deviations of the 109Cd efficiency averaged for all the detectors with respect to their mean value in the [1–8] keV energy region, we obtain a standard deviation of 0.13 , while it increases to 0.3 by considering the detectors independently. These values are better or of the same order that those reported for DAMA/LIBRA-phase2 efficiencies in that energy region (0.3 ) [3]. It can be observed in Figure 8 [25] that 252Cf and 109Cd efficiencies are compatible. In the current analysis, the average of the efficiencies in the ROI derived from the seven 252Cf calibrations will be used. However, according to the evolution of the 109Cd efficiency, constant efficiencies have been assumed for all modules except D0, D4 and D5. During the first year of data collection, modules D4 and D5 showed instabilities until the operation HV was reduced. Consequently, a different (but still constant) efficiency is used in the analysis for D4 during the first 361 days and for D5 during the first 153 days. In the case of D0, one of the PMTs suffered from a gain loss in the last months of the sixth year, and then, a different (but constant) efficiency is used in the analysis for the last 200 days of that year.
@@
After selecting bulk scintillation events by the cut on the BDT parameter, the event rate is evaluated in one-day binning using the corresponding live time. High-rate periods which are more than 3 standard deviations over the annually averaged detection rate below 3 keV are removed (see Ref. [14] for more details). The time corresponding to the rejected periods is discounted from the effective live time used for calculating the effective exposure. The events arriving within 1 s from a cosmic muon triggering the veto system had been removed before applying the BDT cut and the live time had also been conveniently corrected [10].
Results and Discussion
@@
where represents the number of events in the ROI in the time bin for detector , obtained by correcting the measured event count using the live time for that specific temporal bin and detector, along with the corresponding acceptance efficiency; is the Poisson uncertainty associated with the event count, also corrected by the corresponding live time and efficiency; and denotes the expected number of events in that particular time bin and detector, including a hypothetical dark matter component.
is expected to diminish over time because there are background contributions from radioactive isotopes with half-lives on the order of a few years, primarily 210Pb (T1/2=22.3 yr), 3H (T1/2=12.3 yr) and 22Na (T1/2=2.6 yr). For detectors D6, D7 and D8, contributions from Te and I, cosmogenically produced isotopes with shorter half-lives, are also relevant. Accurately modeling this background rate decrease is crucial to avoid biasing the fit. Our background modeling is based on the independent determination of the contamination levels of the crystals and other detector components by different techniques, followed by the Monte Carlo (MC) simulation within the Geant4 package [27]. This background modeling is used to describe the background evolution in time for every detector.
Then, we model as:
where is the probability distribution function sampled from the MC model, describing the background evolution at time bin for detector ; is the mass of every module; and and represent energy and time intervals, respectively. and are free parameters for each detector, while represents the DM annual modulation amplitude. It is set to 0 to test the null hypothesis and allowed to vary freely for the modulation hypothesis. is a time-invariant component of the measured rate that includes both the background contributions produced by isotopes with very long half-lives and a noise contribution not explained by the background model (related to the excess below 3 keV previously commented) but found at a constant rate in the data, as well as the average component of a hypothetical contribution from DM interactions.
In the fit, the period and the phase are fixed at one year and to June 2, respectively, in order to directly compare ANAIS–112 with DAMA/LIBRA results, as they appear in Ref. [3]. We perform two independent fits: in the energy region [2–6] keV, which can be compared with the results from the total accumulated exposure of DAMA/NaI and DAMA/LIBRA, and in the [1–6] keV region, which can be compared with those of DAMA/LIBRA-phase2. Results of the fit in the [1–6] keV energy region are shown in Figure 1. Table 2 shows the results of the fit in the different energy regions analyzed for ANAIS–112 6 years data (this work), together with those of COSINE–100 full dataset [23], and DAMA/LIBRA [3]. It can be observed that the ANAIS–112 results for 6-year exposure are compatible with the absence of modulation within one standard deviation and incompatible with DAMA/LIBRA at 4.0 and 3.5 C.L. for [1–6] and [2–6] keV energy regions, respectively 222These values take into account the uncertainty in the DAMA/LIBRA result, unlike our previous publications. The improvement in sensitivity of our result makes more relevant the contribution from the DAMA/LIBRA uncertainty.. More information is available in the Supplemental Material [25]: fit results corresponding to the [1–6] keV, [2–6] keV energy regions with the individual p-values and /ndf (Figures 10 and 11), residuals after subtracting the non-modulated component for each detector with the DAMA/LIBRA modulation signal superimposed (Figure 13, and a table with the results obtained for all the nuisance parameters considered in the fits, Table 4).
We assess our sensitivity to the DAMA/LIBRA signal as the ratio , which directly gives in units the C.L. at which we can test the DAMA/LIBRA result. Then, the standard deviations for the modulation amplitude obtained in the best fit, cpd/ton/keV for both [1–6] keV and [2–6] keV, correspond to sensitivities of (4.20.4) and (4.10.3), respectively, where the uncertainty corresponds to the 68% C.L. DAMA/LIBRA result uncertainty. This is a highly significant result, showing strong inconsistency with the DAMA/LIBRA signal at an unprecedented level of statistical significance (4 ). In the following, we have studied how some systematics could affect such a direct comparison between ANAIS–112 and DAMA/LIBRA results.
We have carried out 20 000 toy MC simulations of experiments equivalent to 6 years of ANAIS–112 data (taking into account both the time-dependent background evolution and the measured efficiencies), with and without adding the modulation observed by the DAMA/LIBRA experiment. Table 3 shows the bias of the fitting procedure estimated from this analysis for both the modulation and null hypotheses, besides the standard deviation obtained from the fits. No bias is observed and similar standard deviation than found in the ANAIS–112 6 years results are obtained, as shown in Table 2.
| Energy region | Experiment | Exposure | p-value | |
| (kgy) | mod hyp | (cpd/ton/keV) | ||
| [1–6] keV | ANAIS–112 (this work) | 625.75 | 0.156 | -0.42.5 |
| COSINE–100 (full dataset) | 358.00 | — | 1.72.9 | |
| DAMA/LIBRA-phase2 | 1126.40 | 0.513 | 10.51.1 | |
| [2–6] keV | ANAIS–112 (this work) | 625.75 | 0.596 | 1.12.5 |
| COSINE–100 (full dataset) | 358.00 | — | 5.33.1 | |
| DAMA/NaI+DAMA/LIBRA | 2462.09 | 0.935 | 10.20.8 | |
| (cpd/ton/3.3 keV) | ||||
| [6.7–20] keV | ANAIS–112 (this work) | 625.75 | 0.566 | 0.02.3 |
| COSINE–100 (full dataset) | 358.00 | — | 1.32.7 | |
| DAMA/NaI+DAMA/LIBRA | 2462.09 | 0.935 | 10.20.8 |
Then, we have introduced variations in the detection efficiencies of different magnitudes around the mean value and analyzed the modulation hypothesis. The modulation amplitudes obtained by fitting the simulated rate evolution show distributions centered in the DAMA/LIBRA value, with a standard deviation that increases with the fluctuation in the efficiencies introduced, as expected. Figure 2 (upper panel) shows the distribution of the modulation amplitudes derived from the fits in the [1–6] keV energy region for fluctuations in the efficiency of 0, 2, 5 and 10, while the ANAIS–112 result is shown as a dashed line. The measured efficiency variations in ANAIS–112 are on the order of 0.3 , which implies a negligible probability of obtaining a null result when a modulation as large as that observed by DAMA/LIBRA is present in the data. It is worth highlighting that variations as large as 5 are required to yield a non-negligible probability of hindering a true modulation in the data. On the other hand, the value of the fit is highly sensitive to the efficiency fluctuations, as shown in the lower panel of Figure 2. The values obtained by ANAIS are strongly incompatible with fluctuations larger than 2 .
| Energy region | Bias [null hyp] | [null hyp] | Bias [DAMA ] | [DAMA ] |
| (cpd/ton/keV) | (cpd/ton/keV) | (cpd/ton/keV) | (cpd/ton/keV) | |
| [1–6] keV | 0.010.02 | 2.340.01 | 0.000.02 | 2.320.01 |
| [2–6] keV | 0.000.02 | 2.490.01 | 0.010.02 | 2.500.01 |
| [6.7–20] keV | 0.000.02 | 2.230.01 | 0.010.02 | 2.250.01 |
According to the distribution of the modulation amplitude values, /ndf = 23.87/5 (p-value = 2.30) is obtained for the DAMA/LIBRA hypothesis, taking into account DAMA/LIBRA result uncertainty, and /ndf = 3.10/5 (p-value = 0.685) for the null hypothesis, both corresponding to single-hit events in the [1–6] keV energy region. Similar values are obtained in the [2–6] keV energy region. For events with multiplicity 2, we obtain /ndf = 2.26/5 (p-value = 0.812) in the [1–6] keV energy region, in full agreement with the results for single-hits.
Figure 4 shows the modulation amplitude on the scale of sodium nuclear recoils, for 3.3 keV energy bins, for single-hit events, combining the data from the nine ANAIS modules (black dots). DAMA/LIBRA-phase2 (blue triangles) and COSINE–100 data (gray diamonds) are also shown for comparison. This figure assumes constant QF for sodium: 0.2 in the case of ANAIS–112 detectors and 0.3 in the case of DAMA/LIBRA. Because the relation between QF for sodium and iodine in DAMA/LIBRA and ANAIS–112 is the same, the energy scale of this figure can be directly converted into iodine nuclear recoils for both datasets, but not for COSINE–100 data. The ANAIS modulation amplitudes derived within this assumption in the [6.7–20] keV correspond to /ndf = 4.22/4 (p-value = 0.376) for the null hypothesis, and /ndf = 22.98/4 (p-value = 1.28) for the DAMA/LIBRA modulation hypothesis, supporting again with high statistical significance the incompatibility between both experimental results.
Finally, Figure 5 displays in dark blue line the ANAIS–112 sensitivity projection in the [1–6] keV energy region following Ref. [31], conveniently updated to the effective exposure, background level and detection efficiency presented in this work. Similar sensitivities are obtained for the [2–6] keV energy region. Cyan band takes into account the 68% uncertainty in . The black dot is the sensitivity derived from the 6-year result presented here, in good concordance with our estimate. This result supports our expectation of achieving a 5 sensitivity to the DAMA/LIBRA result by the end of 2025.
Summarizing, 6-year exposure ANAIS–112 results are incompatible with DAMA/LIBRA at 4.0 and 3.5 C.L. for [1–6] and [2–6] keV energy regions and at 4.2 C.L. in the [6.7–20] ([22.2–66.7]) keV region assuming a constant QF for sodium (iodine). According to our sensitivity prospects, the scheduled ANAIS–112 full dataset (completed by the end of 2025) will be able to provide a robust refutation of the DAMA/LIBRA signal at 5 level for ER and for NR, even assuming different scintillation QF for DAMA/LIBRA and ANAIS–112 crystals [29, 30]. However, it is worth to remark that the comparison in terms of NR is affected by relevant systematics related with the uncertainties in the knowledge of the scintillation quenching factors for sodium and iodine recoils. The non-proportionality of the NaI(Tl) light yield in the ROI relevant for the annual modulation analysis may partially contribute to these uncertainties. Therefore, a better modeling of both, the quenching factors and the non-proportionality of NaI(Tl) is needed. We are working to improve the estimates of the QF in ANAIS–112 crystals using onsite 252Cf calibrations, but a better evaluation of the QF for DAMA/LIBRA crystals is also necessary.
Acknowledgements
Author contributions statement
The main analysis presented in this article was conducted by I.C. supervised by M.M and M.L.S. The manuscript was written by I.C, M.M. and M.L.S. All authors have read and agreed to the published version of the manuscript. Authors are listed alphabetically by their last names. E.G., A.O.d.S. and M.L.S. contributed to the design of the experiment; J.A., S.C., A.O.d.S., and M.L.S. contributed to the setting-up and commissioning of the experiment; J.A., J.Ap., S.C., D.C., I.C., M.M., Y.O., A.O.d.S., T.P., C.S. and M.L.S. performed calibration and maintenance of the experiment; M.M. developed the DAQ software tools; I.C., M.M. and M.L.S. developed analysis tools; I.C., E.G., M.M., J.P. and M.L.S. analyzed the data; S.C. and T.P. developed background simulation codes; E.G., A.O.d.S. and J.P. performed radiopurity measurements; I.C., E.G., M.M. and J.P. contributed to sensitivity estimates.
Data and code availability
The data and scripts used for preparing the figures and reproducing the results presented in this article will be available at the website of the ORIGINS Excellence Cluster: https://www.origins-cluster.de/odsl/dark-matter-data-center/available-datasets/anais.
Supplemental Material
Detector stability
Figure 8 shows the time evolution of the total detection efficiency in the ROI of ANAIS–112 ([1–6] keV), estimated using both, 109Cd calibration events (in blue) and neutron calibration events (in orange). Neutron calibrations are not evenly distributed along the data taking, but are concentrated in the second half. The solid lines correspond to the mean values for each detector, and the shaded regions represent the standard deviations. Both are compatible within uncertainties. 109Cd data allow to consider constant efficiencies in all the modules, but D0, D4 and D5, as explained in the introduction section.
Low energy background
@@ Acknowledgements
The results of the fits for the annual modulation analysis in the three energy regions considered for the nine modules are shown in Figures 10, 11 and 12. /ndf and p-values under the modulation hypothesis are also individually displayed for each module, besides the global analysis /ndf and p-values for the null and modulation hypothesis and the corresponding to the best fit for the latter. In all the energy regions and detectors, we obtain good fits to the PDF built with our MC background model time evolution. We can conclude that our data are fully consistent with no modulation and incompatible with DAMA/LIBRA. While Table 2 summarizes the fit results for the modulation amplitude and corresponding p-value, Table 4 reports on the nuisance parameters derived from the fits, both background index and for each module, which measures the deviation of the fit from the MC background model detector by detector. Only detector 5 shows systematically values of below 85%. Figure 13 shows the results of the fit after subtracting the non-modulated component. The modulation observed by DAMA/LIBRA is shown in the same plot for comparison.

### v2

@@ Abstract
While astronomical and cosmological evidence about dark matter (DM) accumulates, its nature remains elusive [1]. Over the past thirty years, various excesses above background have been reported in both indirect and direct dark matter searches, most of which have been resolved. However, a few of them persist and deserve further investigation [2]. One of the most puzzling results is that from the DAMA/LIBRA experiment, taking data at the Laboratori Nazionali del Gran Sasso (Italy), which reports a positive detection for more than twenty years [3, 4]. DAMA/LIBRA detectors observe an annual modulation in the detection rate at low energies, below 6 keV (in electron-equivalent energy 111In this Letter, we will use keV for referring to electron-equivalent energies while nuclear recoils energies will be expressed in keV), consistent with the predictions of weakly interacting massive particles (WIMPs) distributed in the galactic halo [6, 7], and having cross-sections with nucleons in the range of cm2 [3]. The statistical significance of such a detection is overwhelming, but the origin of such a modulation has not yet been determined. No other experiment has observed a compatible signal [8, 9], but the uncertainties and unknowns in both the dark matter particle and halo models make it very difficult to disprove the interpretation of the result in terms of a DM signal in a model-independent way [2]. This is the objective of ongoing or recently decommissioned experiments, such as ANAIS–112 [10, 11, 12, 13, 14] and COSINE–100 [15, 16], or other upcoming projects like COSINE-100U [17], COSINUS [18], SABRE [19, 20, 21], and PICOLON [22].
@@
The DAMA/LIBRA experiment is expected to have completed the data taking by the end of 2024. The final results should then be released. The COSINE–100 experiment was in operation from 2016 to 2023 in the YangYang Underground Laboratory in Korea. After concluding data collection, the COSINE–100 collaboration recently released results from the full dataset [23], corresponding to 6.4 years of operation and a total effective exposure of 358 kgyr. The ANAIS–112 experiment, in operation since August 2017 at the Canfranc Underground Laboratory (LSC) in Spain, consists of 112.5 kg of NaI(Tl), distributed in nine detectors, labeled in the following D0 to D8, and produced by the same provider as those used in the COSINE–100 experiment (Alpha Spectra Inc.). A blank module, similar in design to the other modules but without the NaI(Tl) crystal, operates in the same experimental space within an independent shielding, but integrated into the same data acquisition system, to monitor non-NaI(Tl) scintillation events. The trigger threshold is set-up at photoelectron level in each PMT and a coincidence within 200 ns between the two PMT signals of the same module is required for triggering that module. Multiple-hit events are defined as those where two or more module-triggers occur within a coincidence window of 1 s, while for single-hit events only one module triggers in that window. In ANAIS, the muon vetoes are readout by an independent DAQ system and they do not contribute neither to the definition of multiple-hit events nor to the dead time. Details on the ANAIS–112 experimental set-up, detector performance and previous results can be found elsewhere [10, 11, 12, 13, 24, 14]. ANAIS–112 will complete more than 8 years of data by the end of 2025, achieving a sensitivity exceeding the 5 confidence level according to our prospects (see Figure 9 within the Supplemental Material [25]). Here, we present the analysis of the annual modulation corresponding to six years of ANAIS–112 data for a total effective exposure of 625.75 kgyr, which confirms our sensitivity estimates.
@@
First, we briefly describe the key features of the analysis pipeline applied to the 6 years data, which in some aspects differs from that applied to obtain previous results [11, 12, 13, 14].
@@
A new filtering procedure to remove non-bulk scintillation events is applied, following Ref. [24, 26], as done in Ref. [14] for the reanalysis of the 3-year exposure. This filtering protocol is based on Boosted Decision Trees (BDT) and the training does not use background events at all, but uses events from 252Cf calibrations in the [1–2] keV energy region resulting mainly from nuclear elastic scattering from neutrons as signal and non-NaI(Tl) bulk scintillation events from the blank module in an equivalent energy range, as noise. The efficiency of this filtering procedure is derived both from 252Cf and 109Cd calibrations data, and their respective time evolution in the [1–6] keV energy region along the data taking is shown in the Figure 3 of the Supplemental Material [25]. By considering the deviations of the 109Cd efficiency averaged for all the detectors with respect to their mean value in the [1–8] keV energy region, we obtain a standard deviation of 0.13 , while it increases to 0.3 by considering the detectors independently. These values are better or of the same order that those reported for DAMA/LIBRA-phase2 efficiencies in that energy region (0.3 ) [3]. It can be observed [25] that 252Cf and 109Cd efficiencies are compatible. In the current analysis, the average of the efficiencies in the ROI derived from the seven 252Cf calibrations will be used. However, according to the evolution of the 109Cd efficiency, constant efficiencies have been assumed for all modules except D0, D4 and D5. During the first year of data collection, modules D4 and D5 showed instabilities until the operation HV was reduced. Consequently, a different (but still constant) efficiency is used in the analysis for D4 during the first 361 days and for D5 during the first 153 days. In the case of D0, one of the PMTs suffered from a gain loss in the last months of the sixth year, and then, a different (but constant) efficiency is used in the analysis for the last 200 days of that year.
@@
After selecting bulk scintillation events by the cut on the BDT parameter, the event rate is evaluated in one-day binning using the corresponding live time. High-rate periods which are more than 3 standard deviations over the annually averaged detection rate below 3 keV are removed (see Ref. [14] for more details) 222The effect of this cut on the annual modulation analysis presented later in this manuscript was checked with a toy MC. By injecting a modulation with the DAMA/LIBRA amplitude over a background sampled with ANAIS–112 features and applying a cut on periods having a rate at ±3 sigma level of the average, the bias observed in the modulation amplitude recovered after the analysis is (0.00 0.02) cpd/ton/keV..
@@
The time corresponding to the rejected periods is discounted from the effective live time used for calculating the effective exposure. The events arriving within 1 s from a cosmic muon triggering the veto system had been removed before applying the BDT cut and the live time had also been conveniently corrected [10].
The accumulated exposure used for the annual modulation analysis corresponding to the six years of data of the ANAIS–112 experiment is summarized in Table 1 within the Supplemental Material [25]. It also details the dead time, down time, percentage of live time rejected in the analysis, and the corresponding effective exposure after subtracting the latter.
where represents the number of events in the ROI in the time bin for detector , obtained by correcting the measured event count using the live time for that specific temporal bin and detector, along with the corresponding acceptance efficiency; is the Poisson uncertainty associated with the event count, also corrected by the corresponding live time and efficiency, whose uncertainties are combined in quadrature; and denotes the expected number of events in that particular time bin and detector, including a hypothetical dark matter component.
is expected to diminish over time because there are background contributions from radioactive isotopes with half-lives on the order of a few years, primarily 210Pb (T1/2=22.3 yr), 3H (T1/2=12.3 yr) and 22Na (T1/2=2.6 yr). For detectors D6, D7 and D8, contributions from Te and I, cosmogenically produced isotopes with shorter half-lives, are also relevant. Accurately modeling this background rate decrease is crucial to avoid biasing the fit. Our background modeling is based on the independent determination of the contamination levels of the crystals and other detector components by different techniques 333Crystal and other detector components contamination levels introduced in our background modeling have been determined by using different techniques. For the most relevant crystal contaminations we used coincidence measurements (for 40K, 22Na and other cosmogenically produced isotopes), alpha rate determination (for 210Pb-210Bi-210Po), and Bi-Po sequences (for 238U and 232Th chains). For other detector components, HPGe spectroscopy was used, being the PMTs the most relevant background contribution. A thorough description can be found in [27]., followed by the Monte Carlo (MC) simulation within the Geant4 package [27]. This background modeling is used to describe the background evolution in time for every detector.
Then, we model as:
where is the probability distribution function sampled from the MC model, describing the background evolution at time bin for detector ; is a constant probability distribution function that accounts for the noise contribution not explained by the background model (related to the excess below 3 keV previously commented) but found at a constant rate in the data, and the average component of a hypothetical contribution from DM interactions; is the mass of every module; and and represent energy and time intervals, respectively. and are nuisance parameters in the fitting procedure: represents the background index in the considered energy region, while measures the relative weight of the MC estimated time dependence with respect to the total rate. represents the DM annual modulation amplitude, and and are the angular frequency and the phase of the modulation searched for in the data. For the standard halo model, they correspond to a period of 1 year and June, 2, respectively. is set to 0 to test the null hypothesis and allowed to vary freely for the modulation hypothesis.
| Energy region | Experiment | Exposure (kgy) | p-value mod hyp | (cpd/ton/keV) |
| [1–6] keV | ANAIS | 625.75 | 0.156 | -0.42.5 |
| COSINE | 358.00 | 1.72.9 | ||
| DAMA | 1126.40 | 0.513 | 10.51.1 | |
| [2–6] keV | ANAIS | 625.75 | 0.596 | 1.12.5 |
| COSINE | 358.00 | 5.33.1 | ||
| DAMA | 2462.09 | 0.935 | 10.20.8 | |
| (cpd/ton/3.3 keV) | ||||
| [6.7–20] keV | ANAIS | 625.75 | 0.566 | 0.02.3 |
| COSINE | 358.00 | 1.32.7 | ||
| DAMA | 2462.09 | 0.935 | 10.20.8 |
| Energy region | Bias [null hyp] | [null hyp] | Bias [DAMA ] | [DAMA ] |
| (cpd/ton/keV) | (cpd/ton/keV) | (cpd/ton/keV) | (cpd/ton/keV) | |
| [1–6] keV | 0.010.02 | 2.340.01 | 0.000.02 | 2.320.01 |
| [2–6] keV | 0.000.02 | 2.490.01 | 0.010.02 | 2.500.01 |
| [6.7–20] keV | 0.000.02 | 2.230.01 | 0.010.02 | 2.250.01 |
In the fit, the period and the phase are fixed at one year and to June 2, respectively, in order to directly compare ANAIS–112 with DAMA/LIBRA results, as they appear in Ref. [3]. We perform two independent fits: in the energy region [2–6] keV, which can be compared with the results from the total accumulated exposure of DAMA/NaI and DAMA/LIBRA, and in the [1–6] keV region, which can be compared with those of DAMA/LIBRA-phase2. Results of the fit in the [1–6] keV energy region are shown in Figure 1. Table 1 shows the results of the fit in the different energy regions analyzed for ANAIS–112 six-years data (this work), together with those of COSINE–100 full dataset [23], and DAMA/LIBRA [3]. It can be observed that the ANAIS–112 results for 6-year exposure are compatible with the absence of modulation within one standard deviation and incompatible with DAMA/LIBRA at 4.0 and 3.5 C.L. for [1–6] and [2–6] keV energy regions, respectively 444These values take into account the uncertainty in the DAMA/LIBRA result, unlike our previous publications. The improvement in sensitivity of our result makes more relevant the contribution from the DAMA/LIBRA uncertainty.. More information is available in the Supplemental Material [25]: fit results corresponding to the [1–6] keV, [2–6] keV energy regions with the individual p-values and /ndf (Figures 5 and 6), residuals after subtracting the non-modulated component for each detector with the DAMA/LIBRA modulation signal superimposed (Figure 8), and a table with the results obtained for all the nuisance parameters considered in the fits, Table 2).
We assess our sensitivity to the DAMA/LIBRA signal as the ratio , which directly gives in units the C.L. at which we can test the DAMA/LIBRA result. Then, the standard deviations for the modulation amplitude obtained in the best fit, cpd/ton/keV for both [1–6] keV and [2–6] keV, correspond to sensitivities of (4.20.4) and (4.10.3), respectively, where the uncertainty corresponds to the 68% C.L. DAMA/LIBRA result uncertainty. Alternatively, if we define our sensitivity as , where is the quadrature combination of both uncertainties, we obtain statistical significance of 3.8 and 3.9 for the [1–6] and [2–6] keV energy regions, respectively.
This is a highly significant result, showing strong inconsistency with the DAMA/LIBRA signal at an unprecedented level of statistical significance.
We have studied how some systematics could affect such a direct comparison between ANAIS–112 and DAMA/LIBRA results using toy MC simulations of experiments equivalent to 6 years of ANAIS–112 data, with and without adding the modulation observed by the DAMA/LIBRA experiment. Table 2 shows the bias of the fitting procedure estimated from this analysis for both the modulation and null hypotheses, besides the standard deviation obtained from the fits. No bias is observed and similar standard deviations than found in the ANAIS–112 six-years results, as shown in Table 1, are obtained. We have also considered the effect of introducing additional fluctuation in the efficiency, as described in more detail in the End Matter section, finding negligible results in all cases.
According to the distribution of the modulation amplitude values, /ndf = 24.03/5 (p-value = 2.14) is obtained for the DAMA/LIBRA hypothesis, taking into account DAMA/LIBRA result uncertainty, and /ndf = 3.14/5 (p-value = 0.678) for the null hypothesis, both corresponding to single-hit events in the [1–6] keV energy region. Similar values are obtained in the [2–6] keV energy region. For events with multiplicity 2, we obtain /ndf = 2.26/5 (p-value = 0.812) in the [1–6] keV energy region, in full agreement with the results for single-hits.
Figure 3 shows the modulation amplitude on the scale of sodium nuclear recoils, for 3.3 keV energy bins, for single-hit events, combining the data from the nine ANAIS modules (black dots). DAMA/LIBRA-phase2 (blue triangles) and COSINE–100 data (gray diamonds) are also shown for comparison. This figure assumes constant QF for sodium: 0.2 in the case of ANAIS–112 detectors and 0.3 in the case of DAMA/LIBRA. Because the relation between QF for sodium and iodine in DAMA/LIBRA and ANAIS–112 is the same, the energy scale of this figure can be directly converted into iodine nuclear recoils for both datasets, but not for COSINE–100 data. The ANAIS modulation amplitudes derived within this assumption in the [6.7–20] keV correspond to /ndf = 4.22/4 (p-value = 0.376) for the null hypothesis, and /ndf = 22.98/4 (p-value = 1.28) for the DAMA/LIBRA modulation hypothesis, supporting again with high statistical significance the incompatibility between both experimental results. It is worth highlighting that Figure 3 is not a simple re-scaling of the energy axis from Figure 2.
Summarizing, 6-year exposure ANAIS–112 results are incompatible with DAMA/LIBRA at 4.0 and 3.5 C.L. for [1–6] and [2–6] keV energy regions and at 4.2 C.L. in the [6.7–20] ([22.2–66.7]) keV region assuming a constant QF for sodium (iodine), while compatible with full dataset COSINE-100 results, as shown in Figure 6. According to our sensitivity prospects (see Figure 9 within the Supplemental Material), the scheduled ANAIS–112 full dataset (completed by the end of 2025) will be able to provide a robust refutation of the DAMA/LIBRA signal at 5 level for ER and for NR, even assuming different scintillation QF for DAMA/LIBRA and ANAIS–112 crystals [31, 32]. However, it is worth to remark that the comparison in terms of NR is affected by relevant systematics related with the uncertainties in the knowledge of the scintillation quenching factors for sodium and iodine recoils. The non-proportionality of the NaI(Tl) light yield in the ROI relevant for the annual modulation analysis may partially contribute to these uncertainties. Therefore, a better modeling of both, the quenching factors and the non-proportionality of NaI(Tl) is needed. We are working to improve the estimates of the QF in ANAIS–112 crystals using onsite 252Cf calibrations, but a better evaluation of the QF for DAMA/LIBRA crystals is also necessary.
Acknowledgements
Author contributions statement
The main analysis presented in this article was conducted by IC supervised by MM and MLS. The manuscript was written by IC, MM and MLS. All authors have read and agreed to the published version, and are listed alphabetically by their last names. EG, AOdS and MLS contributed to the experiment design; JA, SC, AOdS and MLS contributed to the experiment setting-up and commissioning; JA, JAp, SC, DC, IC, MM, YO, AOdS, TP, CS and MLS performed calibration and maintenance tasks; MM developed DAQ software tools; IC, MM and MLS developed analysis tools; IC, EG, MM, JP and MLS analyzed the data; SC and TP developed background simulations; EG, AOdS and JP performed radiopurity measurements; IC, EG, MM and JP contributed to sensitivity estimates.
Data and code availability
The data that support the findings of this Letter are openly available at the website of the ORIGINS Excellence Cluster: https://www.origins-cluster.de/odsl/dark-matter-data-center/available-datasets/anais.
Supplemental Material
Detector stability
Figure 3 shows the time evolution of the total detection efficiency in the ROI of ANAIS–112 ([1–6] keV), estimated using both, 109Cd calibration events (in blue) and neutron calibration events (in orange). Neutron calibrations are not evenly distributed along the data taking, but are concentrated in the second half. The solid blue and orange lines correspond to the mean values for each detector, and the shaded regions represent the standard deviations. Both are compatible within uncertainties. 109Cd data allow to consider constant efficiencies in all the modules, but D0, D4 and D5, as explained in the introduction section. The efficiency used in the analysis is also represented in Figure 3 by a solid magenta line.
Low energy background
Exposure
@@ Acknowledgements
The results of the fits for the annual modulation analysis in the three energy regions considered for the nine modules are shown in Figures 5, 6 and 7. /ndf and p-values under the modulation hypothesis are also individually displayed for each module, besides the global analysis /ndf and p-values for the null and modulation hypothesis and the corresponding to the best fit for the latter. In all the energy regions and detectors, we obtain good fits to the PDF built with our MC background model time evolution. We can conclude that our data are fully consistent with no modulation and incompatible with DAMA/LIBRA. Fit results for the modulation amplitude and corresponding p-values were shown in Table-I in the Letter. Table 2 reports on the nuisance parameters derived from the fits, both background index and for each module, which measures the deviation of the fit from the MC background model detector by detector. Only detector 5 shows systematically values of below 85%. Figure 8 shows the results of the fit after subtracting the non-modulated component. The modulation observed by DAMA/LIBRA is shown in the same plot for comparison.
@@
ANAIS Sensitivity prospects
Figure 9 displays in dark blue line the ANAIS–112 sensitivity projection in the [1–6] keV energy region following Ref. [33], conveniently updated to the effective exposure, background level and detection efficiency presented in this work. Similar sensitivities are obtained for the [2–6] keV energy region. Cyan band takes into account the 68% uncertainty in . The black dot is the sensitivity derived from the 6-year result presented here, in good concordance with our estimate. This result supports our expectation of achieving a 5 sensitivity to the DAMA/LIBRA result by the end of 2025.